OpenStudy (anonymous):

Find the four arithmetic means between 3 and 28. A. 7, 12, 17, 22 B. 8, 14, 18, 24 C. 9, 14, 19, 24 D. 8, 13, 18, 23

OpenStudy (amistre64):

divide the range between them into 5 equal parts ....

OpenStudy (texaschic101):

that pretty much summed it up right there

OpenStudy (anonymous):

To find arithmetic mean between any 2 numbers , p and q Use the formula \[\huge \frac{ p-q }{ n+1 }\] to find the common difference between the means. The common difference here is \[\huge \frac{ 28-3 }{ 4+1 }=5 \] Adding 5 to the first term gives you the first arithmetic mean, adding 5 to that gives you the second mean, etc. Therefore, the arithmetic means are 8, 13, 18, and 23.

OpenStudy (anonymous):

3+5 =8 8+5=13 13+5=18 18+5=23

OpenStudy (anonymous):

8,13,18,23 are the arithmatic means means

OpenStudy (anonymous):

okay so if i had to find the four arithmetic means between 12 and -13. now would i do that? @No.name

OpenStudy (anonymous):

Yes use the method i gave you , it works everytime

OpenStudy (anonymous):

okay so i just do \[\frac{ -13-12 }{ 4+1 }\]

OpenStudy (anonymous):

Wait OS is lagging a bit i will response wait

OpenStudy (anonymous):

Yes you can do that, Ask your teacher or professor for confirmation

OpenStudy (anonymous):

okay when i did that i got \[\frac{ -25 }{ 5 }\] @No.name